### v3

@@ 4.1 Closed-set
Closed-set questions refer to a type of question where the possible answer options are predefined and limited to a finite set. The evaluation is usually performed on task-specific datasets. In this case, the responses can be naturally judged by benchmark metrics. For example, InstructBLIP [51] reports the accuracy on ScienceQA [100], as well as the CIDEr score [101] on NoCaps [102]. The evaluation settings are typically zero-shot [51, 84] or finetuning [51, 29]. The first setting often selects a wide range of datasets covering different general tasks and splits them into held-in and held-out datasets. After tuning on the former, zero-shot performance is evaluated on the latter with unseen datasets or even unseen tasks. In contrast, the second setting is often observed in the evaluation of domain-specific tasks. For example, LLaVA [16] reports finetuned performance on ScienceQA [100]. LLaVA-Med [29] reports results on biomedical VQA [103].
The above evaluation methods are usually limited to a small range of selected tasks or datasets, lacking a comprehensive quantitative comparison. To this end, some efforts have endeavored to develop new benchmarks specially designed for MLLMs [104, 105]. For example, Fu et al. [104] construct a comprehensive evaluation benchmark MME that includes a total of 14 perception and cognition tasks. All instruction-answer pairs in MME are manually designed to avoid data leakage. MMBench [105] is a benchmark specifically designed for evaluating multiple dimensions of model capabilities, using ChatGPT to match open responses with pre-defined choices. Video-ChatGPT [87] and Video-Bench [106] focus on video domains and propose specialized benchmarks as well as evaluation tools for assessment.
@@ 4.2 Open-set
In contrast to the closed-set questions, the responses to open-set questions can be more flexible, where MLLMs usually play a chatbot role. Because the content of the chat can be arbitrary, it would be trickier to judge than the closed-ended output. The criterion can be classified into manual scoring, GPT scoring, and case study. Manual scoring requires humans to assess the generated responses. This kind of approach often involves hand-crafted questions that are designed to assess specific dimensions. For example, mPLUG-Owl [107] collects a visually related evaluation set to judge capabilities like natural image, diagram, and flowchart understanding. Similarly, GPT4Tools [90] builds two sets for the finetuning and zero-shot performance, respectively, and evaluates the responses in terms of thought, action, arguments, and the whole.
@@
Since manual assessment is labor intensive, some researchers have explored rating with GPT, namely GPT scoring. This approach is often used to evaluate performance on multimodal dialogue. LLaVA [16] proposes to score the responses via text-only GPT-4 in terms of different aspects, such as helpfulness and accuracy. Specifically, 30 images are sampled from the COCO [108] validation set, each associated with a short question, a detailed question, and a complex reasoning question via self-instruction on GPT-4. The answers generated by both the model and GPT-4 are sent to GPT-4 for comparison. Subsequent works follow this idea and prompt ChatGPT or GPT-4 to rate results [29] or judge which one is better [109].
@@
A main issue of applying text-only GPT-4 for evaluation is the judge is only based on translated text content, such as captions or bounding box coordinates, without accessing the image [29]. Thus, it may be questionable to set GPT-4 as the performance upper bound in this case. With the release of the vision interface of GPT, some works exploit the more advanced GPT-4V model to assess the performance of MLLMs. For example, Woodpecker [63] adopts the GPT-4V model to judge the response quality of model answers. The evaluation is expected to be more accurate than using text-only GPT-4 since GPT-4V has direct access to the image.
@@
Since the benchmark evaluation is not comprehensive enough, a supplementary approach is to compare the different capabilities of MLLMs through case studies. For instance, some studies evaluate two typical advanced commercial-use models, GPT-4V and Gemini. Yang et al. [110] perform in-depth qualitative analysis on GPT-4V by crafting a series of samples across various domains and tasks, spanning from preliminary skills, such as caption and object counting, to complex tasks that require world knowledge and reasoning, such as joke understanding and indoor navigation as an embodied agent. Wen et al. [111] make a more focused evaluation of GPT-4V by designing samples targeting automatic driving scenarios. Fu et al. [112] carry out a comprehensive evaluation on Gemini-Pro by comparing the model against GPT-4V. The results suggest that GPT-4V and Gemini exhibit comparable visual reasoning abilities in spite of different response styles.
@@ 5 EXTENSIONS
Granularity Support. To facilitate better interaction between agents and users, researchers have developed MLLMs with finer support of granularities in terms of model inputs and outputs. On the input side, models that support finer control from user prompts are developed progressively, evolving from image to region [24] and even pixels [25]. Specifically, Shikra [24] supports region-level input and understanding. Users may interact with the assistant more flexibly by referring to specific regions, which are represented in bounding boxes of natural language forms. Ferret [113] takes a step further and supports more flexible referring by devising a hybrid representation scheme. The model supports different forms of prompts, including point, box, and sketch. Similarly, Osprey [25] supports point input by utilizing a segmentation model [10]. Aided by the exceptional capabilities of the pre-trained segmentation model, Osprey enables specifying a single entity or part of it with a single click. On the output side, grounding capabilities are improved in line with the development of input support. Shikra [24] supports response grounded in the image with box annotations, resulting in higher precision and finer referring experience. LISA [114] further supports mask-level understanding and reasoning, which makes pixel-level grounding possible.
Modality Support. Increased support for modalities is a tendency for MLLM studies. On the one hand, researchers have explored adapting MLLMs to support the input of more multimodal content, such as 3D point cloud [115]. On the other hand, MLLMs are also extended to generate responses of more modalities, such as image [116], audio [117], and video [118]. For example, NExT-GPT [119] proposes a framework that supports inputs and outputs of mixed modalities, specifically, combinations of text, image, audio, and video, with the help of diffusion models [120] attached to the MLLM. The framework applies an encoder-decoder architecture and puts LLM as a pivot for understanding and reasoning.
@@
A typical tendency is to adapt MLLMs to more specific real-life scenarios. For example, some works develop agents that interact with the real world, e.g. user-friendly assistants specially designed for Graphical User Interface (GUI), as exemplified by CogAgent [32], AppAgent [122], and Mobile-Agent [123]. Researchers also develop embodied agents [31, 19] that can perform reasoning, navigation, and manipulation in the real world, facilitating the development of automatic agents that can execute tasks for humans. In general, these assistants excel in planning and performing each step to fulfill tasks specified by users, acting as helpful agents for humans.
Another line is to augment MLLMs with specific skills for solving tasks in different domains, e.g. document understanding [30] and medical domains [29]. For document understanding, mPLUG-DocOwl [124] utilizes various forms of document-level data for tuning, resulting in an enhanced model in OCR-free document understanding. TextMonkey [30] incorporates multiple tasks related to document understanding to improve model performance. Similarly, MLLMs can also be trained to accommodate traditional vision tasks such as visual grounding [125, 126]. Compared with traditional methods [127, 13], MLLMs unify the I/O format and streamline the whole learning and inference process. Specifically, it is feasible to recast the grounding task into a conditioned box coordinate prediction task under a unified language modeling objective [24, 52, 28]. The model is trained to predict the coordinates of specified objects in the form of natural language. MLLMs can also be extended to medical domains by instilling specialized knowledge. For example, LLaVA-Med [29] develops assistants specialized in medical image understanding and question answering by injecting domain knowledge.
Efficient MLLMs. Recently, using lightweight MLLMs for efficient deployment has gained increased popularity [128, 129, 130]. These models are meticulously designed and optimized for more economical utilization or resource-limited scenarios without compromising too much on model performance.
From a model perspective, various techniques have been explored to facilitate efficient training and inference. For instance, MobileVLM [54] explores developing small-size variants of MLLMs for resource-limited scenarios. Some designs and techniques are utilized for deployment on mobile devices, such as LLMs of smaller size and quantization techniques to speed up computation. Similarly, MiniCPM-V [129] builds efficient MLLMs for end-side computation. A Q-Former [28] is adopted to cut down the number of visual tokens for each patch of the image.
From a data perspective, Bunny [130] comprehensively investigates efficient data selection and combination schemes for model training. The obtained models achieve performance on par with MLLMs of larger parameter sizes.
@@ 6 MULTIMODAL HALLUCINATION
Multimodal hallucination refers to the phenomenon of responses generated by MLLMs being inconsistent with the image content [63]. The fundamental problem has received increased attention. In this section, we briefly introduce related concepts and research development.
@@ 6.1 Preliminaries
Multimodal hallucinations can be categorized into three types [131]:
@@
    Existence Hallucination is a common type, meaning that models incorrectly decide the existence of objects.
@@
    Attribute Hallucination means falsely describing the attributes of certain objects, e.g. failure to identify a dog’s color.
@@
    Relationship Hallucination is a more complex type of hallucination. It refers to false descriptions of relationships between objects, such as relative positions.
In what follows, we first introduce evaluation methods, which are useful to gauge the performance of methods for mitigating hallucinations. Then, we discuss mitigation methods of different kinds of approaches.
@@ 6.2 Evaluation methods
CHAIR [132] is an early metric that evaluates hallucination levels in open-ended captions. The metric measures the proportion of sentences with hallucinated objects or the proportion of hallucinated objects in all the objects mentioned. In contrast, POPE [133] is a method that evaluates closed-set choices. Specifically, multiple prompts with binary choices are formulated, each querying if a specific object exists in the image. With a similar evaluation approach, MME [104] provides a more comprehensive evaluation, covering aspects of existence, count, position, and color, as exemplified in [63].
Different from previous approaches that use matching mechanisms to detect and decide hallucinations, some works explore automatic evaluation of text responses via models. For example, HaELM [134] proposes using LLMs as a judge to decide whether MLLMs’ captions are correct against reference captions. In view of the fact that text-only LLMs can only access limited image context and require reference annotations, Woodpecker [63] uses GPT-4V to directly assess model responses grounded in the image.
@@ 6.3 Mitigation methods
According to high-level ideas for mitigating hallucinations, current methods can be roughly divided into three categories: pre-correction, in-process-correction, and post-correction.
Pre-correction. An intuitive solution for hallucination is to collect specialized data (e.g. negative data) and use the data for fine-tuning, thus achieving models with fewer hallucinations.
LRV-Instruction [135] introduces a visual instruction tuning dataset to encourage faithful generation. Similarly, LLaVA-RLHF [96] collects human-preference pairs and finetunes models with reinforcement learning techniques.
In-process-correction. Another line is to make improvements in architectural design or feature representation. These works try to explore the reasons for hallucinations and design remedies to mitigate them in the generation process. For example, HallE-Switch [131] introduces a continuous controlling factor to control the extent of imagination in model output during inference.
Post-correction. Different from previous paradigms, post-correction mitigates hallucinations in a post-remedy way. For example, Woodpecker [63] is a training-free framework for hallucination correction. Specifically, the method incorporates expert models to supplement contextual information of the image and crafts a pipeline to correct hallucinations step by step.
@@ 7.1 Multimodal in-context learning
ICL is one of the important emergent abilities of LLMs. The essence of the technique is prompting the model with a few examples as guidance to make it easier for the model to answer the query. There are two good traits of ICL: (1) The crux of ICL is to learn from analogy [136], thus largely reducing the requirement of data samples. (2) ICL is usually implemented in a training-free way [136] and can be flexibly integrated into various frameworks at inference time.
@@
Table 9: A simplified example of the template to structure an M-ICL query, adapted from [81]. For illustration, we list two in-context examples and a query divided by a dashed line. {instruction} and {response} are texts from the data sample. <image> is a placeholder to represent the multimodal input (an image in this case). <BOS> and <EOS> are tokens denoting the start and the end of the input to the LLM, respectively.
In the context of MLLM, ICL has been extended to more modalities, leading to Multimodal ICL (M-ICL). At inference time, M-ICL can be implemented by adding a demonstration set, i.e. a set of in-context samples, to the original sample. In this case, the template can be extended as illustrated in Table 9.

### v4

@@ 4.1 Closed-set
Closed-set questions refer to a type of question where the possible answer options are predefined and limited to a finite set. The evaluation is usually performed on task-specific datasets. In this case, the responses can be naturally judged by benchmark metrics [60, 20, 70, 102, 76, 103, 104, 101]. For example, InstructBLIP [60] reports the accuracy on ScienceQA [116], as well as the CIDEr score [117] on NoCaps [118] and Flickr30K [119]. The evaluation settings are typically zero-shot [60, 102, 104, 105] or finetuning [60, 20, 70, 76, 103, 101, 35, 105]. The first setting often selects a wide range of datasets covering different general tasks and splits them into held-in and held-out datasets. After tuning on the former, zero-shot performance is evaluated on the latter with unseen datasets or even unseen tasks. In contrast, the second setting is often observed in the evaluation of domain-specific tasks. For example, LLaVA [20] and LLaMA-Adapter [76] report finetuned performance on ScienceQA [116]. LLaVA-Med [35] reports results on biomedical VQA [120, 121, 122].
The above evaluation methods are usually limited to a small range of selected tasks or datasets, lacking a comprehensive quantitative comparison. To this end, some efforts have endeavored to develop new benchmarks specially designed for MLLMs [123, 124, 125, 126, 127, 128, 129]. For example, Fu et al. [123] construct a comprehensive evaluation benchmark MME that includes a total of 14 perception and cognition tasks. All instruction-answer pairs in MME are manually designed to avoid data leakage. MMBench [124] is a benchmark specifically designed for evaluating multiple dimensions of model capabilities, using ChatGPT to match open responses with pre-defined choices. Video-ChatGPT [130] and Video-Bench [131] focus on video domains and propose specialized benchmarks as well as evaluation tools for assessment. There are also evaluation strategies designed to evaluate a specific aspect of the model [102], as exemplified by POPE [132] for assessment of hallucination degree.
@@ 4.2 Open-set
In contrast to the closed-set questions, the responses to open-set questions can be more flexible, where MLLMs usually play a chatbot role. Because the content of the chat can be arbitrary, it would be trickier to judge than the closed-ended output. The criterion can be classified into manual scoring, GPT scoring, and case study. Manual scoring requires humans to assess the generated responses. This kind of approach often involves hand-crafted questions that are designed to assess specific dimensions. For example, mPLUG-Owl [81] collects a visually related evaluation set to judge capabilities like natural image understanding, diagram, and flowchart understanding. Similarly, GPT4Tools [107] builds two sets for the finetuning and zero-shot performance, respectively, and evaluates the responses in terms of thought, action, arguments, and the whole.
@@
Since manual assessment is labor intensive, some researchers have explored rating with GPT, namely GPT scoring. This approach is often used to evaluate performance on multimodal dialogue. LLaVA [20] proposes to score the responses via text-only GPT-4 in terms of different aspects, such as helpfulness and accuracy. Specifically, 30 images are sampled from the COCO [133] validation set, each associated with a short question, a detailed question, and a complex reasoning question via self-instruction on GPT-4. The answers generated by both the model and GPT-4 are sent to GPT-4 for comparison. Subsequent works follow this idea and prompt ChatGPT [81] or GPT-4 [70, 104, 101, 35, 105] to rate results [81, 70, 104, 101, 35] or judge which one is better [103].
@@
A main issue of applying text-only GPT-4 as an evaluator is that the judge is only based on image-related text content, such as captions or bounding box coordinates, without accessing the image [35]. Thus, it may be questionable to set GPT-4 as the performance upper bound in this case. With the release of the vision interface of GPT, some works [77, 134] exploit a more advanced GPT-4V model to assess the performance of MLLMs. For example, Woodpecker [77] adopts GPT-4V to judge the response quality of model answers based on the image. The evaluation is expected to be more accurate than using text-only GPT-4 since GPT-4V has direct access to the image.
@@
A supplementary approach is to compare the different capabilities of MLLMs through case studies. For instance, some studies evaluate two typical advanced commercial-use models, GPT-4V and Gemini. Yang et al. [135] perform in-depth qualitative analysis on GPT-4V by crafting a series of samples across various domains and tasks, spanning from preliminary skills, such as caption and object counting, to complex tasks that require world knowledge and reasoning, such as joke understanding and indoor navigation as an embodied agent. Wen et al. [136] make a more focused evaluation of GPT-4V by designing samples targeting automatic driving scenarios. Fu et al. [137] carry out a comprehensive evaluation on Gemini-Pro by comparing the model against GPT-4V. The results suggest that GPT-4V and Gemini exhibit comparable visual reasoning abilities in spite of different response styles.
@@ 5 Extensions
Granularity Support. To facilitate better interaction between agents and users, researchers have developed MLLMs with finer support of granularities in terms of model inputs and outputs. On the input side, models that support finer control from user prompts are developed progressively, evolving from image to region [28, 138, 139] and even pixels [29, 140, 141]. Specifically, Shikra [28] supports region-level input and understanding. Users may interact with the assistant more flexibly by referring to specific regions, which are represented in bounding boxes of natural language forms. Ferret [141] takes a step further and supports more flexible referring by devising a hybrid representation scheme. The model supports different forms of prompts, including point, box, and sketch. Similarly, Osprey [29] supports point input by utilizing a segmentation model [9]. Aided by the exceptional capabilities of the pre-trained segmentation model, Osprey enables specifying a single entity or part of it with a single click. On the output side, grounding capabilities are improved in line with the development of input support. Shikra [28] supports response grounded in the image with box annotations, resulting in higher precision and finer referring experience. LISA [142] further supports mask-level understanding and reasoning, which makes pixel-level grounding possible.
Modality Support. Increased support for modalities is a tendency for MLLM studies. On the one hand, researchers have explored adapting MLLMs to support the input of more multimodal content, such as 3D point cloud [143, 144, 41, 145]. On the other hand, MLLMs are also extended to generate responses of more modalities, such as image [146, 147, 32, 148], audio [149, 150, 147, 32], and video [32, 151]. For example, NExT-GPT [32] proposes a framework that supports inputs and outputs of mixed modalities, specifically, combinations of text, image, audio, and video, with the help of diffusion models [152, 153] attached to the MLLM. The framework applies an encoder-decoder architecture and puts LLM as a pivot for understanding and reasoning.
@@
A typical tendency is to adapt MLLMs to more specific real-life scenarios. MobileVLM [63] explores developing small-size variants of MLLMs for resource-limited scenarios. Some designs and techniques are utilized for deployment on mobile devices, such as LLMs of smaller size and quantization techniques to speed up computation. Other works develop agents that interact with real-world [154, 41, 155], e.g. user-friendly assistants specially designed for Graphical User Interface (GUI), as exemplified by CogAgent [44], AppAgent [43], and Mobile-Agent [45]. These assistants excel in planning and guiding through each step to fulfill a task specified by users, acting as helpful agents for human-machine interaction. Another line is to augment MLLMs with specific skills for solving tasks in different domains, e.g. document understanding [38, 39, 156, 157] and medical domains [35, 37, 36]. For document understanding, mPLUG-DocOwl [38] utilizes various forms of document-level data for tuning, resulting in an enhanced model in OCR-free document understanding. TextMonkey [39] incorporates multiple tasks related to document understanding to improve model performance. Apart from conventional document image and scene text datasets, position-related tasks are added to reduce hallucinations and help models learn to ground responses in the visual information. MLLMs can also be extended to medical domains by instilling knowledge of the medical domain. For example, LLaVA-Med [158] injects medical knowledge into vanilla LLaVA [20] and develops an assistant specialized in medical image understanding and question answering.
@@ 6 Multimodal Hallucination
Multimodal hallucination refers to the phenomenon of responses generated by MLLMs being inconsistent with the image content [77]. As a fundamental and important problem, the issue has received increased attention. In this section, we briefly introduce some related concepts and research development.
@@ 6.1 Preliminaries
Current research on multimodal hallucinations can be further categorized into three types [159]:
@@
    Existence Hallucination is the most basic form, meaning that models incorrectly claim the existence of certain objects in the image.
@@
    Attribute Hallucination means describing the attributes of certain objects in a wrong way, e.g. failure to identify a dog’s color correctly. It is typically associated with existence hallucination since descriptions of the attributes should be grounded in objects present in the image.
@@
    Relationship Hallucination is a more complex type and is also based on the existence of objects. It refers to false descriptions of relationships between objects, such as relative positions and interactions.
In what follows, we first introduce some specific evaluation methods (§6.2), which are useful to gauge the performance of methods for mitigating hallucinations (§6.3). Then, we will discuss in detail the current methods for reducing hallucinations, according to the main categories each method falls into.
@@ 6.2 Evaluation Methods
CHAIR [160] is an early metric that evaluates hallucination levels in open-ended captions. The metric measures the proportion of sentences with hallucinated objects or hallucinated objects in all the objects mentioned. In contrast, POPE [132] is a method that evaluates closed-set choices. Specifically, multiple prompts with binary choices are formulated, each querying if a specific object exists in the image. The method also covers more challenging settings to evaluate the robustness of MLLMs, with data statistics taken into consideration. The final evaluation uses a simple watchword mechanism, i.e. by detecting keywords “yes/no”, to convert open-ended responses into closed-set binary choices. With a similar evaluation approach, MME [123] provides a more comprehensive evaluation, covering aspects of existence, count, position and color, as exemplified in [77].
Different from previous approaches that use matching mechanisms to detect and decide hallucinations, HaELM [161] proposes using text-only LLMs as a judge to automatically decide whether MLLMs’ captions are correct against reference captions. In light of the fact that text-only LLMs can only access limited image context and require reference annotations, Woodpecker [77] uses GPT-4V to directly assess model responses grounded in the image. FaithScore [162] is a more fine-grained metric based on a routine that breaks down descriptive sub-sentences and evaluates each sub-sentence separately. Based on previous studies, AMBER [163] is an LLM-free benchmark that encompasses both discriminative tasks and generative tasks and involves three types of possible hallucinations (see §6.1).
@@ 6.3 Mitigation Methods
According to high-level ideas, the current methods can be roughly divided into three categories: pre-correction, in-process-correction, and post-correction.
Pre-correction. An intuitive and straightforward solution for hallucination is to collect specialized data (e.g. negative data) and use the data for fine-tuning, thus resulting in models with fewer hallucinated responses.
LRV-Instruction [164] introduces a visual instruction tuning dataset. Apart from common positive instructions, the dataset incorporates delicately designed negative instructions at different semantic levels to encourage responses faithful to the image content. LLaVA-RLHF [112] collects human-preference pairs and finetunes models with reinforcement learning techniques, leading to models more aligned with less hallucinated answers.
In-process-correction. Another line is to make improvements in architectural design or feature representation. These works try to explore the reasons for hallucinations and design corresponding remedies to mitigate them in the generation process.
HallE-Switch [159] performs an empirical analysis of possible factors of object existence hallucinations and hypothesizes that existence hallucinations derive from objects not grounded by visual encoders, and they are actually inferred based on knowledge embedded in the LLM. Based on the assumption, a continuous controlling factor and corresponding training scheme are introduced to control the extent of imagination in model output during inference.
VCD [165] suggests that object hallucinations derive from two primary causes, i.e. statistical bias in training corpus and strong language prior embedded in LLMs. The authors take notice of the phenomenon that when injecting noise into the image, MLLMs tend to lean towards language prior rather than the image content for response generation, leading to hallucinations. Correspondingly, this work designs an amplify-then-contrast decoding scheme to offset the false bias.
HACL [166] investigates the embedding space of vision and language. Based on the observation, a contrastive learning scheme is devised to pull paired cross-modal representation closer while pushing away non-hallucinated and hallucinated text representation.
Post-correction. Different from previous paradigms, post-correction mitigates hallucinations in a post-remedy way and corrects hallucinations after output generation. Woodpecker [77] is a training-free general framework for hallucination correction. Specifically, the method incorporates expert models to supplement contextual information of the image and crafts a pipeline to correct hallucinations step by step. The method is interpretable in that intermediate results of each step can be checked, and objects are grounded in the image. The other method LURE [167] trains a specialized revisor to mask objects with high uncertainty in the descriptions and regenerates the responses again.
@@ 7.1 Multimodal In-Context Learning
ICL is one of the important emergent abilities of LLMs. There are two good traits of ICL: (1) Different from traditional supervised learning paradigms that learn implicit patterns from abundant data, the crux of ICL is to learn from analogy [168]. Specifically, in the ICL setting, LLMs learn from a few examples along with an optional instruction and extrapolate to new questions, thereby solving complex and unseen tasks in a few-shot manner [169, 22, 170]. (2) ICL is usually implemented in a training-free manner [168] and thus can be flexibly integrated into different frameworks at the inference stage. A closely related technique to ICL is instruction-tuning (see §3.2), which is shown empirically to enhance the ICL ability [19].
@@
TABLE IX: A simplified example of the template to structure an M-ICL query, adapted from [98]. For illustration, we list two in-context examples and a query divided by a dashed line. {instruction} and {response} are texts from the data sample. <image> is a placeholder to represent the multimodal input (an image in this case). <BOS> and <EOS> are tokens denoting the start and the end of the input to the LLM, respectively.
In the context of MLLM, ICL has been extended to more modalities, leading to Multimodal ICL (M-ICL). Building upon the setting in (§3.2), at inference time, M-ICL can be implemented by adding a demonstration set, i.e. a set of in-context samples, to the original sample. In this case, the template can be extended as illustrated in Table IX. Note that we list two in-context examples for illustration, but the number and the ordering of examples can be flexibly adjusted. In fact, models are commonly sensitive to the arrangement of demonstrations [168, 171].